OpenStudy (anonymous):

-2(v-7)-6v=8(4-v)-3 Help?! i keep getting 15, is that correct?

OpenStudy (anonymous):

-2v + 14 - 6v = 32 - 8v - 3 -8v + 14 = 29 - 8v 14 does not equal to 29 Something wrong with your question

OpenStudy (anonymous):

well once your in the seconsp wouldnt you ad 8v to each side?

OpenStudy (anonymous):

*second

OpenStudy (anonymous):

*second

OpenStudy (anonymous):

yea so you get 14 = 29

OpenStudy (anonymous):

-2(v-7)-6v=8(4-v)-3 -2v+14-6v=32-8v-3 -6v-2v+14=32-3-8v -8v+14=29-8v add 8v to both sidesss.... 14=29? Not true.....

OpenStudy (anonymous):

how did you figure out the problem?